# FragmentGPT: A Unified GPT Model for Fragment Growing, Linking, and Merging in Molecular Design

### v1

@@ 4.2 Main results
Link Link-INVENT (Guo et al., 2023)
@@
Merge Link-INVENT (Guo et al., 2023)
@@ 4.2.1 Fragment Growing
Figure 3 illustrates examples of fragment growing, where the original molecule was obtained from an NMR-based screening (G. Ferreira and D. Andricopulo, 2017). This example highlights the ability of FragmentGPT to effectively transform simple fragments into chemically plausible and drug-like molecules.
@@ 4.2.2 Fragment Linking
dBET6, as visualized in figure 4, is a PROTAC molecule with high structural selectivity that bridges the E3 ubiquitin ligase CRBN and the target protein BRD4 (Winter et al., 2017). We use this example to investigate whether our model can identify reasonable linkage sites and propose a suitable linker, and results are represented in figure 5.
@@ 4.2.3 Fragment Merging
To assess our model’s ability to merge different fragments, we utilized the COVID Moonshot dataset (Boby et al., 2023)—a fully open science, crowd-sourced, and structure-enabled drug discovery campaign against the severe acute respiratory syndrome coronavirus 2 (SARS-CoV-2) main protease (Mpro). This dataset provides an ideal testbed for evaluating fragment-based drug design approaches due to its comprehensive collection of crystallographic fragment hits and subsequent rational design submissions. The submission TRY-UNI-714a760b represents a class of rationally designed molecules that were constructed based on the strategic combination of five distinct fragments identified through crystallographic screening. These fragments were carefully selected and merged by human medicinal chemists who recognized their spatial complementarity and potential for synergistic binding interactions within the Mpro active site. To evaluate our model’s fragment merging capabilities, we challenged it to reproduce this human-designed approach by providing the same five fragments as input and tasking it with generating novel molecular combinations. Remarkably, our model successfully generated four molecules among the twenty four molecules present in the submission series, as shown in figure 6. Our model can efficiently identify a variety of drug-related functional groups, including pyridine rings, amide bonds, and others. This achievement demonstrates our model’s sophisticated ability to identify mergeable sites and recognize the structural compatibility between different fragment scaffolds.
@@ A.1 Pre-training and finetuning dataset
For finetuning, we employed 1 million compounds from the ZINC15 dataset, each docked to the RTCB protein (PDB ID: 4DWQ) implicated in human cancer, as provided by the Cancer dataset of Liu et al. [2023a], and applied them consistently across all baselines. In each finetuning epoch, a random subset of the desired number of molecules was sampled.
@@ A.3 Surrogate model
The surrogate model [Vasan et al., 2023] adopts a simplified BERT-style transformer architecture, widely used in natural language processing. In this framework, tokenized SMILES strings are embedded with positional encodings and passed through a stack of five transformer blocks. Each block consists of a multi-head attention mechanism with 21 heads, followed by dropout, layer normalization with residual connections, and a feedforward subnetwork. The feedforward subnetwork is composed of two dense layers, again followed by dropout and layer normalization with residual connections. After the transformer stack, a final feedforward layer produces the predicted docking score. The validation performance achieved $r^{2}$ values of 0.73 on the RTCB dataset.
@@ A.5 Hyperparameters and architectures
For experimentation, we use a set of 1,280 molecules from the RTCB dataset, selecting those with docking scores in the range of [14,6][-14,-6], following the range reported in [Xuefeng et al., 2024].

### v2

@@ 4.2 Main results
Link Link-INVENT [11]
@@
Merge Link-INVENT [11]
@@ 4.2.1 Fragment Growing
Figure 3 illustrates examples of fragment growing, where the original molecule was obtained from an NMR-based screening [8]. This example highlights the ability of FragmentGPT to effectively transform simple fragments into chemically plausible and drug-like molecules.
@@ 4.2.2 Fragment Linking
dBET6, as visualized in figure 4, is a PROTAC molecule with high structural selectivity that bridges the E3 ubiquitin ligase CRBN and the target protein BRD4 [37]. We use this example to investigate whether our model can identify reasonable linkage sites and propose a suitable linker, and results are represented in figure 5.
@@ 4.2.3 Fragment Merging
To assess our model’s ability to merge different fragments, we utilized the COVID Moonshot dataset [3]—a fully open science, crowd-sourced, and structure-enabled drug discovery campaign against the severe acute respiratory syndrome coronavirus 2 (SARS-CoV-2) main protease (Mpro). This dataset provides an ideal testbed for evaluating fragment-based drug design approaches due to its comprehensive collection of crystallographic fragment hits and subsequent rational design submissions. The submission TRY-UNI-714a760b represents a class of rationally designed molecules that were constructed based on the strategic combination of five distinct fragments identified through crystallographic screening. These fragments were carefully selected and merged by human medicinal chemists who recognized their spatial complementarity and potential for synergistic binding interactions within the Mpro active site. To evaluate our model’s fragment merging capabilities, we challenged it to reproduce this human-designed approach by providing the same five fragments as input and tasking it with generating novel molecular combinations. Remarkably, our model successfully generated four molecules among the twenty four molecules present in the submission series, as shown in figure 6. Our model can efficiently identify a variety of drug-related functional groups, including pyridine rings, amide bonds, and others. This achievement demonstrates our model’s sophisticated ability to identify mergeable sites and recognize the structural compatibility between different fragment scaffolds.
@@ A.1 Pre-training and finetuning dataset
For finetuning, we employed 1 million compounds from the ZINC15 dataset, each docked to the RTCB protein (PDB ID: 4DWQ) implicated in human cancer, as provided by the Cancer dataset of Liu et al. [23], and applied them consistently across all baselines. In each finetuning epoch, a random subset of the desired number of molecules was sampled.
@@ A.3 Surrogate model
The surrogate model [33] adopts a simplified BERT-style transformer architecture, widely used in natural language processing. In this framework, tokenized SMILES strings are embedded with positional encodings and passed through a stack of five transformer blocks. Each block consists of a multi-head attention mechanism with 21 heads, followed by dropout, layer normalization with residual connections, and a feedforward subnetwork. The feedforward subnetwork is composed of two dense layers, again followed by dropout and layer normalization with residual connections. After the transformer stack, a final feedforward layer produces the predicted docking score. The validation performance achieved $r^{2}$ values of 0.73 on the RTCB dataset.
@@ A.5 Hyperparameters and architectures
For experimentation, we use a set of 1,280 molecules from the RTCB dataset, selecting those with docking scores in the range of [14,6][-14,-6], following the range reported in [40].